Gordon Rosel went to his bank to find out how long it will take for $1,800 to amount to $2,310 at 7.75% simple interest.

Please solve Gordon’s problem.

Answer by stanbon(75887)

Gordon Rosel went to his bank to find out how long it will take for $1,800 to amount to $2,310 at 7.75% simple interest.
A(t) = P(1+rt)
2310 = 1800(1+0.0775*t)
1.2833 = 1+0.0775t
0.0775t = 0.2833
t = 3.66 years
---------------------
Cheers,
Stan H.
===============